A factory kept increasing its output by the same percentage. Find the percentage increase if its known output is doubled in the last two years?

Grade:10

1 Answers

Ashwin Muralidharan IIT Madras
12 years ago

Say X was output in the yr before last two years.

 

So nex year the output = X(1+P)

And the year after that output = X(1+P)(1+P), which is given to be twice the output 2 yrs back.

So (1+p)2=2 or 1+P = root(2) = 1.414

So P=0.414.

 

Or percentage change every year = 41.4%.

 

All the best.

Regards,

Ashwin (IIT Madras).